OpenStudy (loser66):

I have to use Wronskian to find out the particular solution of ODE, but I cannot get the answer, please help y"-5y'+6y=2e^t

OpenStudy (loser66):

I got y_1 =e^2t; y_2 =e^3t, and Wronskian is e^5s then ???

OpenStudy (loser66):

plug all of them into the formula, never get the right answer which is e^t. Please, help

OpenStudy (loser66):

one more question: when calculate Wronskian, if I let e^3t first then e^2t, the W(s) = -e^5s. which is different when I let e^2t the e^3t , the W(s) = e^5s. Why??

OpenStudy (anonymous):

Swapping columns of a matrix changes the sign of the determinant; this makes intuitive sense if you look at it as an "oriented" volume of the parallelipid spanned by its columns in regular linear algebra

OpenStudy (loser66):

Oh, yah, so we should be "flexible" in choosing y_1 y_2 to get the W(s) positive, right?

OpenStudy (anonymous):

right; sign is arbitrary and depends on which you call the "first" solution

OpenStudy (loser66):

can someone show me, please

OpenStudy (anonymous):

\[ y_p= - y_1\int \frac{y_2 e^{2t}}{e^{5t}}dt+ y_2\int \frac{y_1 e^{2t}}{e^{5t}}dt \]

OpenStudy (loser66):

yes

OpenStudy (loser66):

there is 2 in numerator since g(s) = 2e^t

OpenStudy (loser66):

the solution in book and in Wolframalpha is e^t. :( !! I didn't get it.

OpenStudy (anonymous):

Oh Sorry, I read \[ g(t) = e^{2t}\]

OpenStudy (anonymous):

Let me fix that

OpenStudy (anonymous):

\[ y_p= - e^{2t}\int \frac{e^{3t} 2 e^{t}}{e^{5t}}dt+ e^{3t}\int \frac{e^{2t} 2 e^{t}}{e^{5t}}dt=\\ -e^{2t}\int2 e^{-t} dt+e^{3t}\int2 e^{-2t}dt=2 e^{t}- e^{t}=e^{t}\\ \]

OpenStudy (loser66):

so, we don't have to plug any arbitrary limits to the integral to find it out? Anyway, thanks a ton... I am stuck with it for 8 hours.

OpenStudy (anonymous):

No, that is the correct way of doing it. Enjoy.

OpenStudy (loser66):

my prof asked us do something like